A silo has the form of a right circular cylinder 20m in diameter and 45m
high. The lateral surface of the silo is to be painted with two coats of
elastomeric paint.

1. If the contract calls for two coats of paint and if it is estimated that
one gallon of paint can cover 30 square meters of surface in one
coating, how many gallons of paint needed to cover the lateral
surface area of the silo?
2. How much will be the contract amount for painting if painting cost is Php 800 per sqm?
1
Expert's answer
2020-11-16T19:17:39-0500

1). Lateral area of silo= 2πhr = πdh = 20×45×π = 2829m2

1 gallon = 30m2

2829m2 = 2829/30 = 94.3 gallons

Since there will be 2 coatings, then total gallons = 2× 94.3 = 188.6gallons


2. ) 1m2 = 800php

2829m2 = 2829 ×800 = 2263200Php